OpenStudy (anonymous):

Let f (x) = -2x+4 and g(x)= 6(x) -7. Find F(x) - g(x).

OpenStudy (anonymous):

What?

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} f (x) = -2x+4&&g(x)= 6(x) -7\\ \qquad \uparrow &&\qquad \uparrow \\ f(x)&-&g(x)\\ \qquad \downarrow &&\qquad \downarrow \\ -2x+4&-&6(x) -7 \end{array}\)